(b) If the rod is a rotating cantilever of 20 cm length with a transverse force F at its free end. Find the maximum force Fmax that will never break the rod due to fatigue.

Clear
Answer #1

The assumed loading configuration is given below:

Assumed diameter = d m

Moment of Inertia

\small I=\frac{\pi d^4}{64}

The max bending moment is(at the base of cantilever)

\small M=FL=0.2F\: Nm

Therefore max bending stress is

\small \sigma=\frac{My}{I}=\frac{0.2F(d/2)}{\pi d^4/64}=\frac{6.4F}{\pi d^3}=\frac{2.0372F}{d^3}\: Pa

For a point located at the base of cantilever at surface of rod , the above stress will vary in sign from positive to negative.

Hence maximum stress is

\small \sigma_{max}=\frac{2.0372F}{d^3}, \sigma_{min}=-\frac{2.0372F}{d^3}\: Pa

Therefore mean stress is

\small \sigma_m=\frac{\sigma_{max}+\sigma_{min}}{2}=0

Stress amplitude is

\small \sigma_a=\frac{\sigma_{max}-\sigma_{min}}{2}=\frac{2.0372F}{d^3}\: Pa

Calculation of endurance limit for infinite cycles(Ref Shigley Chap-6)

Since material is not given we assume AISI 1040 steel(machined)

Tensile strength Sut= 620 MPa

Yield strength Sy= 415 MPa

The rotating beam beam specimen(base) endurance limit is

\small S'_e=0.5S_u_t=0.5(620)=310\: MPa

The surface condition factor is (for machined surface)

\small k_a=aS_u_t^b=(4.51)(620^{-0.265})=0.8207

The size factor is(assuming d <51 mm)

\small k_b=1.24(1000d)^{-0.107}=0.6065d^{-0.107}

Loading factor kc= 1 (for bending)

Temperature factor kd= 1 (assumed room temperature)

Reliability factor ke= 0.897 (assumed 90% reliability)

Hence the actual endurance limit is

\small S_e=k_ak_bk_ck_dk_eS'_e=(0.8207)(0.6065d^{-0.107})(1)(1)(0.897)(310)=138.41d^{-0.107}\: MPa

Applying modified Goodman criterion

\small \frac{\sigma_a}{S_e}+\frac{\sigma_m}{S_u_t}=\frac{1}{n}

n= fatigue factor of safety=1 (for max F)

Hence

\small \frac{2.0372F/d^3}{138.41d^{-0.107}\times10^6}+\frac{0}{620}=\frac{1}{1}

Or

\small 1.4718\times10^{-8}d^{-2.893}F=1

Hence

\small F_{max}=67.94\times10^6d^{2.893}=826\: N

For example if d= 20 mm= 0.02 m

\small F_{max}=67.94\times10^6(0.02^{2.893})=826\: N